# Stability theory over toroidal or Novikov-type base and canonical modifications

### v1

@@ 1. Introduction
In this paper, we introduce more extensive algebraic (partly stack-theoretic) generalities with such perspective, which leads to introduce e.g., isotrivial degenerating families over toroidal bases, Novikov type valuation rings of high rank, and their analytic analogues. In particular, focusing on the context of stability and moduli theories of varieties, we introduce the notion of generalized test configurations and their limit versions, and show that it is useful to treat the following types of degenerations in “irrational direction” in literatures, a more canonical and unified manners:
@@
In our other papers we plan to discuss and explore these one by one, but this paper still focuses on the general framework. In particular, the latter half of this paper gives some generalization of the $\Theta$-stability theory and stable reduction theorem ([HL14] and [AHLH23] respectively). Here, recall $\Theta:=[\mathbb{A}^{1}/\mathbb{G}_{m}]$ over a field $k$ is the quotient stack whose coarse moduli is just one point which is used as a source of map to express equivariant isotrivial families over $\mathbb{A}^{1}$. [HL14, AHLH23] introduces the notion of $\Theta$-strata and $\Theta$-stratification on (highly non-separated) algebraic stacks, as a stacky reminicsents of the Hesselink-Kempf-Kirwan-Ness stratification [Kir84, Nes84] and the Harder-Narasimhan filtration of coherent sheaves [HN74]. We call a natural higher rank version of it as higher $\Theta$-strata. Also, as an irrational analogue of DVR, we introduce Novikov type spectra $\overline{\Delta_{\xi}}(R)$ (Definition 2.3). This gives an extension of the valuation ring $R$.
@@ 2.1. Setup of the bases
Consider the minimum lattice $N^{\prime}\subset N\oplus\mathbb{Z}$ whose spanning $\mathbb{R}$-linear subspace contains $\xi^{\prime}$, which we suppose to have rank $r$. We take the projective system of rational polyhedral cones $\tau^{\prime}$ of $N^{\prime}$ which contains $\xi^{\prime}$. Then consider the natural projection
@@
Note that for a fixed $\xi$ and $\tau$, there is a canonical choice of $c$ such that the lift of $\tau$ as $\tau^{\prime}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ contains $(v_{i},1)$ for each of the basis $\{v_{i}\}_{i}$ of the regular generator of the cone $\tau\cap N$ thanks to the irrationality of $\xi$. $\pi$ is also naturally decided in some context later. In such cases, we abbreviate and denote as $\overline{\Delta_{\xi}}(R,\pi)$ (resp., $\widetilde{\Delta_{\xi}}(R,\pi)$), $\overline{\Delta_{\xi}}(R)$ (resp., $\widetilde{\Delta_{\xi}}(R)$), or even $\overline{\Delta_{\xi}}$ (resp., $\widetilde{\Delta_{\xi}}$) for brevity. These are also regarded as versions of N-spectra.
@@ 2.2. Generalized test configurations and examples
    An affine algebraic test deformation of $(Y,\xi)$ over $U_{\tau}$ is algebraic positive (resp., negative) weight deformation (or $Y$ is algebraic positive (resp., negative) $\xi$-weight deformation) over $U_{\overline{\tau}}$ if
@@
      $\tau\ni\xi$ (resp., $-\tau\ni\xi$).
@@
We define the algebraic subtorus $T\subset\mathbb{G}_{m}^{r}$ as $T:=\mathbb{G}_{m}\otimes N$. We take generators of the defining ideal of $X\subset\mathbb{A}_{\mathbbm{k}}^{l}$, say $f_{1},\cdots f_{a}$, which are even universal Gröbner basis of the generating ideal of $X$ ([BCRV22, §5.1]).
@@
    Suppose $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$ is an algebraic negative weight deformation (Definition 2.6) of $(Y,\xi)$. In particular, $\xi\in-\tau$. Then, its general fiber is a negative $\xi$-weight deformation in the sense of [CH22, Definition 1.7].
@@
For the item (1), we can assume that $\tau=\mathbb{Z}_{\geq 0}^{r}$, $N=\mathbb{Z}^{r}$. Then we approximate $\xi$ by $\xi_{i}=(\frac{r_{i,1}}{m_{1}},\frac{r_{i,2}}{m_{1}},\cdots,\frac{r_{i,r}}{m_{1}})\in-\tau\cap N_{\mathbb{Q}}$, with nonnegative integers $r_{i,j}(j=1,\cdots,r)$ and $m_{i}$s so that a convergence $\xi_{i}\to\xi$ holds for $i\to+\infty$. Then take a morphism $\varphi_{\vec{r}}\colon\mathbb{A}^{1}\to U_{\tau}$ defined by $t\mapsto(t^{r_{i,1}},t^{r_{i,2}},\cdots,t^{r_{i,r}})$ and consider the base change $\mathcal{Y}\times_{p,U_{\tau},\varphi_{\vec{r}}}\mathbb{A}^{1}$, which we denote as $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}_{1}$. It can be naturally seen as an affine test configuration but we modify the $\mathbb{G}_{m}$-action $\sigma^{\prime}_{i}$ on it by compositing with the $m_{i}$-th power map $\mathbb{G}_{m}\to\mathbb{G}_{m}$. In particular, $p^{\prime}_{i}(\sigma^{\prime}_{i}(z,x))=z^{m_{i}}p^{\prime}_{i}(x)$ for $z\in\mathbb{G}_{m}(\mathbbm{k}),x\in W^{\prime}_{i}$. Then, we take the base change of $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}^{1}$ with $\sigma^{\prime}_{i}$ by the $m_{i}$-th power map $q_{i}\colon\mathbb{A}^{1}\to\mathbb{A}^{1}$ whose target is regarded as the base of $p^{\prime}_{i}$. It is straightforward confirmation that the sequence, with respect to $i=1,2,\cdots$, of the obtained fiber product $p_{i}\colon W_{i}(:=W^{\prime}_{i}\times_{p^{\prime}_{i},\mathbb{A}^{1},q_{i}}\mathbb{A}^{1})\to\mathbb{A}^{1}$ with the $\mathbb{G}_{m}$-action induced by $\sigma^{\prime}_{i}$ satisfies the conditions of the [CH22, Definition 1.7], with $c_{i}$ of loc.cit as $m_{i}$, $\mu_{i}$ of loc.cit as $1$, hence the latter being auxiliary parameters. Suppose $Z$ is the maximum Artin scheme supported at the $T$-invariant point $p_{\tau}\in U_{\tau}$, such that $\mathcal{Y}|_{Z}$ is trivial family. Such $Z$ exists due to the Zorn’s lemma combined with the formal smoothness ([Gro64, Part I, Part IV]) of ${\rm Aut}_{T}(Y)$ which is an algebraic group as the stabilizer of $[Y]$ in the multi-Hilbert scheme. By using this $Z$, $k_{i}$ of loc.cit can be written as $m_{i}{\rm length}\mathcal{O}_{\varphi_{\vec{r}}^{*}Z}$, which explains the nature of the $\xi$-weight in loc.cit. ∎
@@ 3.1. Toroidal stack-theoretic setup
Consider the mapping stack ${\rm Map}(\Theta^{r},\mathcal{M})$ of $\Theta^{r}$ to $\mathcal{M}$, as an algebraic stack, after [AHR20, 5.10, 5.11], [AHR19, 6.23]. Extending [HL14, Definition 2.1.1], a higher $\Theta$-stratum of rank $r$ in $\mathcal{M}$ consists of a union of connected components $\mathcal{Z}^{+}\subset{\rm Map}(\Theta^{r},\mathcal{M})$ such that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed embedding. If we fix an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for a rational polyhedral cone $\tau$, then we also call higher $\Theta$-stratum as higher $\Theta$-stratum for the cone $\tau$.
Further, if $\mathcal{M}$ is quotient stack of the form $[M/T]$ where $M$ is an algebraic $\mathbbm{k}$-scheme $M$ and $T$ is an algebraic torus again, we often write ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=[Z^{+}/T]$ for some $T$-invariant closed subscheme $Z^{+}\subset M$.
@@
The following lemma lies the ring-theoretic aspect of Definition 2.3, as both appear in Theorem 3.8 later.
@@
Suppose $(R,v=v_{R})$ is a valuation local ring, such that the associated graded ring $gr_{v}(R):=\oplus_{w\in{\rm Im}(v)}\{a\in R\mid v(a)\geq w\}/\{a\in R\mid v(a)>w\}$ is isomorphic to a semigroup ring over a field for a submonoid of $\mathbb{R}^{r}$ with some $r\in\mathbb{Z}_{>0}$ as a graded ring (which holds e.g. if $R$ is of the form $k[[\Gamma_{\geq 0}]]$ in Definition 2.1), and $\xi,N,M_{\xi\geq 0}$ are as Notation 1. Now, we set
@@
Set $b:=\frac{v(\pi)}{\langle\xi,m\rangle}>0$ and put $\tilde{v}(\sum a_{i}t^{m_{i}}):=\min_{i}\{v(a_{i})+b\langle m_{i},\xi\rangle\}$ for $a_{i}\in R,m_{i}\in M_{\xi\geq 0}$. This $\tilde{v}$ defines a valuation on $R[[M_{\xi\geq 0}]]$ and consider its corresponding filtration $\mathcal{F}_{\tilde{v}}$ as $\mathcal{F}_{\tilde{v}}(x)=\{f\in R[[M_{\xi\geq 0}]]\mid\tilde{v}(f)\geq x\}$ for each $x\in\mathbb{R}$. It induces a filtration $\overline{\mathcal{F}}$ on $R_{\rm new}(\xi)$ as a quotient of $R[[M_{\xi\geq 0}]]$, and the corresponding graded ring is clearly $({\rm gr}_{v}(R))[M_{\xi}/\mathbb{Z}_{\geq 0}m]$ which is still a semigroup ring for a submonoid of $\mathbb{R}^{r}$ for a bigger $r$. Since it is an integral domain, $\overline{\mathcal{F}}$ also induces a valuation on $R_{\rm new}(\xi)$. ∎
@@
Even if $R$ is of the form $k[[\Gamma_{\geq 0}]]$ in Definition 2.1, hence a valuation ring, $R_{\rm new}(\xi)^{\prime}$ is not necessarily of the form $k[[\Gamma_{\geq 0}]]$ for a different $\Gamma_{\geq 0}$ as in Definition 2.1, and it is not a valuation ring in general. For instance, such a pathology often happens if we continue the above process $R\rightsquigarrow R_{\rm new}(\xi)$ for the same irrational $\xi\in M_{\mathbb{R}}$ twice. Nevertheless, if $R$ is of such form $k[[\Gamma_{\geq 0}]]$, $R_{\rm new}(\xi)$ is again of such form $k[[\Gamma_{\geq 0}+bM_{\xi\geq 0}]]$ for $b\in\mathbb{R}_{>0}$ as in the proof of Lemma 3.5.
@@
In the setting of Definition 2.3, a formal completion of the inductive limit of $R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})$ is nothing but $R_{\rm new}(\xi^{\prime})$ defined in the above Lemma 3.5. Hence, $R_{\rm new}(\xi^{\prime}(=(\xi,c))$ can be understood as the completion of a stalk of $\overline{\Delta_{\xi}}(c,\pi,R)$ of Definition 2.3.
@@ 3.2. Higher $\Theta$-stable reduction theorem
Below, we generalize the semistable reduction type theorem in [AHLH23, §6.1], which was for $r=1$ with particular emphasis on canonicity and uniqueness. Even in that $r=1$ case, the following proof (for the quotient stack case) seems to be different from loc.cit and more toric geometric.
Morally speaking, it says that a one parameter family which strictly degenerates into a higher $\Theta$-strata, there is a canonical modification of the family to avoid the higher $\Theta$-strata i.e., with “less degeneration”. This is a version of stable reduction type theorem, which we later refine in Theorem 3.18.
@@
Let $(R,\mathfrak{m},v)$ be a valuation $\mathbbm{k}$-algebra and $\mathcal{M}$ be an algebraic stack over $\mathbbm{k}$ with affine diagonal.
@@
  • replacing $R$ by another valuation $\mathbbm{k}$-algebra with local $\mathbbm{k}$-morphism $R\to R^{\prime}$, so that the valuation of $R^{\prime}$ restricts to that of $R$.
@@
If $\mathcal{M}$ is a moduli stack of varieties and $\overline{f}$ lifts to a morphism from $\tilde{\rm ST}_{R}(\tau^{\prime})$ to a chart $M$ of $\mathcal{M}$, on which $T$ acts, its restrictions to each of the two strata corresponding to $\tau^{\prime}$ and $\tau$ gives a generalized test configurations in the sense of Definition 2.6.
@@
In particular, we obtain a canonical modification $f_{{\rm new},\xi}\colon\mathop{\mathrm{Spec}}\nolimits R_{\rm new}(\xi)\to\mathcal{M}$ which maps the closed point to outside $\mathcal{Z}^{+}$.
@@
As [HL14, Lemma 4.4.6] (cf., also [AHHLR22]) shows, there is an affine $\mathbbm{k}$-scheme $M$, algebraic $\mathbbm{k}$-split torus $T=N\otimes\mathbb{G}_{m}$ of rank $r$ and a smooth surjective representable morphism $[M/T]\to\mathcal{M}$ which induces ${\rm Filt}^{r}([M/T])\to{\rm Filt}^{r}(\mathcal{M})$, at the level of $\mathbb{Z}^{r}$-filtration stack (as algebraic stacks by [AHR20, 5.10, 5.11], [AHR19, 6.23]) for $r={\rm rank}(N)$ in particular. Hence we reduce the proof to the case when $\mathcal{M}$ is a quotient stack of the form $M$ an algebraic scheme over $\mathbbm{k}$ acted by an algebraic $\mathbbm{k}$-split torus $T$ of dimension $r$. We assume it henceforth.
From the structure of higher $\Theta$-strata on $Z^{+}$, as we assume, there is a $T$-equivariant morphism $Z^{+}\times U_{\tau}\to M$, following Notation 1, for some polyhedral cone $\tau$. Since there is also the (compatible) $T$-action on $M$, they induce a $T$-equivariant rational map $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$. From the condition that $Z^{+}$ gives a union of connected components in the mapping stack, it follows that $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$ is indeed not defined in the neighborhood of $(p_{\tau},\mathfrak{m})$, hence we want to resolve the indeterminancy as follows.
@@
We consider $T$-equivariant local coordinates $z_{1},\cdots,z_{N}$ around the image of the $T$-invariant point $p_{\tau}\in U_{\sigma}$ and consider its pullback (as rational functions) to $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)$ and denote them as $\frac{f_{i}}{g_{i}}$ with holomorphic (algebraic) $f_{i}$s and $g_{i}$s with $T$-semiinvariant $g_{i}$s. Consider the ideal $\mathcal{J}:=\langle\{f_{i},g_{i}\mid 1\leq i\leq N\}\rangle$ which can be decomposed as
@@
where $\vec{m}$ runs over elements of $\tau^{\vee}\cap M$, $z^{\vec{m}}$ means $\prod_{i}z_{i}^{m_{i}}$, and $I_{\vec{m}}$ are ideals of $R$. With respect to the natural partial order on $\tau^{\vee}\cap M$ as $\vec{m}\leq\vec{m^{\prime}}:\Leftrightarrow\vec{m^{\prime}}-\vec{m}\in\tau^{\vee}\cap M$, we have $I_{\vec{m}}\subset I_{\vec{m}}$ if $\vec{m}\leq\vec{m^{\prime}}$. If we take a uniformizer of $R$ as $\pi$ and we can write $I_{\vec{m}}$ as $(\pi^{l_{m}})$ since $R$ is a discrete valuation ring, with $l_{m}\in\mathbb{Z}_{\geq 0}$. We have $l_{\vec{m}}\geq l_{\vec{m^{\prime}}}\geq 0$ if $\vec{m}\leq\vec{m^{\prime}}$.

### v2

@@ 1. Introduction
In this paper, we introduce more extensive algebraic (partly stack-theoretic) generalities with such perspective, which leads to introduce e.g., isotrivial degenerating families over toroidal bases, Novikov type valuation rings of high rank, and their analytic analogues. In particular, focusing on the context of stability and moduli theories of varieties, we introduce the notion of generalized test configurations and their limit versions. Then, together with a companion paper, we show that the generalized notion is useful to treat the following types of degenerations in “irrational direction” in literatures, a more canonical and unified manners:
@@
In our other papers we plan to discuss and explore these one by one, but this paper still focuses on the general framework. In particular, the latter half of this paper gives some generalization of the $\Theta$-stability theory and stable reduction theorem ([HL14] and [AHLH23] respectively), as a core theorem in this paper. Here, recall $\Theta:=[\mathbb{A}^{1}/\mathbb{G}_{m}]$ over a field $k$ is the quotient stack whose coarse moduli is just one point which is used as a source of map to express equivariant isotrivial families over $\mathbb{A}^{1}$. [HL14, AHLH23] introduces the notion of $\Theta$-strata and $\Theta$-stratification on (highly non-separated) algebraic stacks, as a stacky reminicsents of the Hesselink-Kempf-Kirwan-Ness stratification [Kir84, Nes84] and the Harder-Narasimhan filtration of coherent sheaves [HN74]. We call a natural higher rank version of it as higher $\Theta$-strata. Also, as an irrational analogue of DVR, we introduce Novikov type spectra $\overline{\Delta_{\xi}}(R)$ (Definition 2.3). This gives an extension of the valuation ring $R$.
@@ 2.1. Setup of the bases
Consider the minimum lattice $N^{\prime}\subset N\oplus\mathbb{Z}$ whose spanning $\mathbb{R}$-linear subspace contains $\xi^{\prime}$, which we suppose to have rank $r$. This assumption naturally arises in our later application in §3.2. We take the projective system of rational polyhedral cones $\tau^{\prime}$ of $N^{\prime}$ which contains $\xi^{\prime}$. Then consider the natural projection
@@
In the setup we later apply above to Theorem 3.8 and Corollary 3.10, there is a canonical choice of $c$ such that the lift of $\tau$ as $\tau^{\prime}\subset N_{\mathbb{R}}\oplus\mathbb{R}$ contains $(v_{i},1)$ for each of the basis $\{v_{i}\}_{i}$ of the regular generator of the cone $\tau\cap N$ thanks to the irrationality of $\xi$. $\pi$ is also naturally decided in some context later. In such cases, we abbreviate and denote as $\overline{\Delta_{\xi}}(R,\pi)$ (resp., $\widetilde{\Delta_{\xi}}(R,\pi)$), $\overline{\Delta_{\xi}}(R)$ (resp., $\widetilde{\Delta_{\xi}}(R)$), or even $\overline{\Delta_{\xi}}$ (resp., $\widetilde{\Delta_{\xi}}$) for brevity. These are also regarded as versions of N-spectra.
@@ 2.2. Generalized test configurations and examples
    An affine algebraic test deformation of $(Y,\xi)$ over $U_{\tau}$ is algebraic positive (resp., negative) weight deformation (or $Y$ is algebraic scale up (resp., scale down) deformation) over $U_{\overline{\tau}}$ if
@@
      $-\tau\ni\xi$ (resp., $\tau\ni\xi$).  [^4] Our sign convention comes from [Pin74, CH22] and matches with differential geometric intuition, while being opposite with some other algebraic literature.
@@
We define the algebraic subtorus $T\subset\mathbb{G}_{m}^{r}$ as $T:=\mathbb{G}_{m}\otimes N$. We take generators of the defining ideal of $X\subset\mathbb{A}_{\mathbbm{k}}^{l}$, say $f_{1},\cdots f_{a}$, which are even universal Gröbner basis of the defining ideal of $X$ ([BCRV22, §5.1]).
@@
    Suppose $p\colon\mathcal{Y}\twoheadrightarrow U_{\tau}$ is an algebraic negative weight deformation (Definition 2.6) of $(Y,\xi)$. In particular, $\xi\in\tau$. Then, its general fiber is a negative $\xi$-weight deformation in the sense of [CH22, Definition 1.7].
@@
For the item (1), we can assume that $\tau=\mathbb{Z}_{\geq 0}^{r}$, $N=\mathbb{Z}^{r}$. Then we approximate $\xi$ by $\xi_{i}=(\frac{r_{i,1}}{m_{1}},\frac{r_{i,2}}{m_{1}},\cdots,\frac{r_{i,r}}{m_{1}})\in\tau\cap N_{\mathbb{Q}}$, with nonnegative integers $r_{i,j}(j=1,\cdots,r)$ and $m_{i}$s so that a convergence $\xi_{i}\to\xi$ holds for $i\to+\infty$. Then take a morphism $\varphi_{\vec{r}}\colon\mathbb{A}^{1}\to U_{\tau}$ defined by $t\mapsto(t^{r_{i,1}},t^{r_{i,2}},\cdots,t^{r_{i,r}})$ and consider the base change $\mathcal{Y}\times_{p,U_{\tau},\varphi_{\vec{r}}}\mathbb{A}^{1}$, which we denote as $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}_{1}$. It can be naturally seen as an affine test configuration but we modify the $\mathbb{G}_{m}$-action $\sigma^{\prime}_{i}$ on it by compositing with the $m_{i}$-th power map $\mathbb{G}_{m}\to\mathbb{G}_{m}$. In particular, $p^{\prime}_{i}(\sigma^{\prime}_{i}(z,x))=z^{m_{i}}p^{\prime}_{i}(x)$ for $z\in\mathbb{G}_{m}(\mathbbm{k}),x\in W^{\prime}_{i}$. Then, we take the base change of $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}^{1}$ with $\sigma^{\prime}_{i}$ by the $m_{i}$-th power map $q_{i}\colon\mathbb{A}^{1}\to\mathbb{A}^{1}$ whose target is regarded as the base of $p^{\prime}_{i}$. It is straightforward confirmation that the sequence, with respect to $i=1,2,\cdots$, of the obtained fiber product $p_{i}\colon W_{i}(:=W^{\prime}_{i}\times_{p^{\prime}_{i},\mathbb{A}^{1},q_{i}}\mathbb{A}^{1})\to\mathbb{A}^{1}$ with the $\mathbb{G}_{m}$-action induced by $\sigma^{\prime}_{i}$ satisfies the conditions of the [CH22, Definition 1.7], with $c_{i}$ of loc.cit as $m_{i}$, $\mu_{i}$ of loc.cit as $1$, hence the latter being auxiliary parameters. Suppose $Z$ is the maximum Artin scheme supported at the $T$-invariant point $p_{\tau}\in U_{\tau}$, such that $\mathcal{Y}|_{Z}$ is trivial family. Such $Z$ exists due to the Zorn’s lemma combined with the formal smoothness ([Gro64, Part I, Part IV]) of ${\rm Aut}_{T}(Y)$ which is an algebraic group as the stabilizer of $[Y]$ in the multi-Hilbert scheme. By using this $Z$, $k_{i}$ of loc.cit can be written as $m_{i}{\rm length}\mathcal{O}_{\varphi_{\vec{r}}^{*}Z}$, which explains the nature of the $\xi$-weight in loc.cit. ∎
@@ 3.1. Toroidal stack-theoretic setup
Consider the mapping stack ${\rm Map}(\Theta^{r},\mathcal{M})$ of $\Theta^{r}$ to $\mathcal{M}$, as an algebraic stack, after [AHR20, 5.10, 5.11], [AHR19, 6.23], [HLP23, 5.1]. Extending [HL14, Definition 2.1.1], a higher $\Theta$-stratum of rank $r$ in $\mathcal{M}$ consists of a union of connected components $\mathcal{Z}^{+}\subset{\rm Map}(\Theta^{r},\mathcal{M})$ such that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed embedding. If we fix an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for a rational polyhedral cone $\tau$, then we also call higher $\Theta$-stratum as higher $\Theta$-stratum for the cone $\tau$.
Further, if $\mathcal{M}$ is quotient stack of the form $[M/G]$ where $M$ is an algebraic $\mathbbm{k}$-scheme $M$ and $G$ is an algebraic $\mathbbm{k}$-group, we often write ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=[Z^{+}/G]$ for some $G$-invariant locally closed subscheme $Z^{+}\subset M$.
@@
The following lemma lies the ring-theoretic aspect of Definition 2.3 and it appears in Theorem 3.8 later.
@@
Suppose $(R,v=v_{R})$ is a valuation local ring, such that the associated graded ring $gr_{v}(R):=\oplus_{w\in{\rm Im}(v)}\{a\in R\mid v(a)\geq w\}/\{a\in R\mid v(a)>w\}$ is isomorphic to a semigroup ring over a field $\mathbbm{k}$ and a submonoid of $\mathbb{R}^{r}$ with some $r\in\mathbb{Z}_{>0}$ as a graded ring (which holds e.g. if $R$ is of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ for a field $\mathbbm{k}$ in Definition 2.1) and $\xi,N,M_{\xi\geq 0}$ are as Notation 1. Now, we set
@@
Set $b:=\frac{v(\pi)}{\langle\xi,m\rangle}>0$ and put $\tilde{v}(\sum a_{i}t^{m_{i}}):=\min_{i}\{v(a_{i})+b\langle m_{i},\xi\rangle\}$ for $a_{i}\in R,m_{i}\in M_{\xi\geq 0}$. This $\tilde{v}$ defines a valuation on $R[[M_{\xi\geq 0}]]$ and consider its corresponding filtration $\mathcal{F}_{\tilde{v}}$ as $\mathcal{F}_{\tilde{v}}(x)=\{f\in R[[M_{\xi\geq 0}]]\mid\tilde{v}(f)\geq x\}$ for each $x\in\mathbb{R}$. It induces a filtration $\overline{\mathcal{F}}$ on $R_{\rm new}(\xi)$ as a quotient of $R[[M_{\xi\geq 0}]]$, and the corresponding graded ring is clearly $({\rm gr}_{v}(R))[M_{\xi}/\mathbb{Z}_{\geq 0}m]$ which is still a semigroup over $\mathbbm{k}$ ring for a submonoid of $\mathbb{R}^{r}$ for a bigger $r$. Since it is an integral domain, $\overline{\mathcal{F}}$ also induces a valuation on $R_{\rm new}(\xi)$. ∎
@@
Even if $R$ is of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ in Definition 2.1, hence a valuation ring, $R_{\rm new}(\xi)^{\prime}$ is not necessarily of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ for a different $\Gamma_{\geq 0}$ as in Definition 2.1, and it is not a valuation ring in general. For instance, such a pathology often happens if we continue the above process $R\rightsquigarrow R_{\rm new}(\xi)$ for the same irrational $\xi\in M_{\mathbb{R}}$ twice. Nevertheless, if $R$ is of such form $\mathbbm{k}[[\Gamma_{\geq 0}]]$, $R_{\rm new}(\xi)$ is again of such form $\mathbbm{k}[[\Gamma_{\geq 0}+bM_{\xi\geq 0}]]$ for $b\in\mathbb{R}_{>0}$ as in the proof of Lemma 3.5.
@@
In the setting of Definition 2.3, a formal completion of the inductive limit of $R[M^{\prime}\cap\tau^{\prime\vee}]/(\pi-\chi^{p(\vec{0},1)})$ is nothing but $R_{\rm new}(\xi^{\prime})$ for $\xi^{\prime}\in N^{\prime}\otimes\mathbb{R}$ with some $m\in M^{\prime}$ defined in the above Lemma 3.5. (The place of is not a typo.) Hence, $R_{\rm new}(\xi^{\prime}(=(\xi,c))$ can be understood as the completion of a stalk of $\overline{\Delta_{\xi}}(c,\pi,R)$ of Definition 2.3.
@@ 3.2. Higher $\Theta$-stable reduction theorem
Below, we generalize the semistable reduction type theorem in [AHLH23, §6.1], which was for $r=1$ case. Morally speaking, it says that a one parameter family which strictly degenerates into a higher $\Theta$-strata, there is a canonical modification of the family to avoid a higher $\Theta$-strata i.e., with “less degeneration”. This is a version of stable reduction type theorem, and later Theorem 3.18 is its iterated version along a certain kind of stratification (higher $\Theta$-stratification cf., Definition 3.17).
@@
Let $\mathbbm{k}$ be a field, $(R,\mathfrak{m},v)$ be an equicharacteristic divisorial valuation ring which is $\mathbbm{k}$-algebra with $R/\mathfrak{m}\simeq\mathbbm{k}$ (divisorial valuation $\mathbbm{k}$-algebra), or more generally a valuation $\mathbbm{k}$-algebra which satisfies Lemma 3.5, and $\mathcal{M}$ be an algebraic stack over $\mathbbm{k}$ with affine diagonal.
@@
  • replacing $R$ by another divisorial valuation $\mathbbm{k}$-algebra with local $\mathbbm{k}$-morphism $R\to R^{\prime}$, so that the valuation of $R^{\prime}$ restricts to that of $R$.
@@
If $\mathcal{M}$ is a moduli stack of varieties and $\overline{f}$ lifts to a morphism from $\widetilde{\rm ST}_{R}(\tau^{\prime})$ to a chart $M$ of $\mathcal{M}$, on which $T$ acts, its restrictions to each of the two strata corresponding to $\tau^{\prime}$ and $\tau$ gives a generalized test configurations in the sense of Definition 2.6.
@@
In particular, we obtain a canonical modification $f_{{\rm new},\xi}\colon\mathop{\mathrm{Spec}}\nolimits R_{\rm new}(\xi)\to\mathcal{M}$ which maps the closed point to outside $\mathcal{Z}^{+}$ (recall Lemma 3.5).
@@
As [HL14, Lemma 4.4.6] (cf., also [AHHLR22]) shows, there is an affine $\mathbbm{k}$-scheme $M$, algebraic $\mathbbm{k}$-split torus $T^{\prime}$ and a smooth surjective representable morphism $[M/T^{\prime}]\to\mathcal{M}$ which induces ${\rm Filt}^{r}([M/T^{\prime}])\to{\rm Filt}^{r}(\mathcal{M})$, at the level of $\mathbb{Z}^{r}$-filtration stack (as algebraic stacks by [AHR20, 5.10, 5.11], [AHR19, 6.23]) for any $r$. Hence, we reduce the proof to the case when $\mathcal{M}$ is a quotient stack of the form $M$ an algebraic scheme over $\mathbbm{k}$ acted by an algebraic $\mathbbm{k}$-split torus $T^{\prime}$ as above. We assume it henceforth.
From our assumption on the structure of higher $\Theta$-strata on $Z^{+}$ and the Białynicki-Birula type theorem [HL14, 1.4.7], there is a $T$-equivariant morphism $Z^{+}\times U_{\tau}\to M$, following Notation 1, for some polyhedral cone $\tau$. Since there is also the (compatible) $T$-action on $M$, they induce a $T$-equivariant rational map $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$. From the condition that $Z^{+}$ gives a union of connected components in the mapping stack, it follows that $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)\dashrightarrow M$ is indeed not defined in the neighborhood of $(p_{\tau},\mathfrak{m})$, hence we want to resolve the indeterminancy as follows.
@@
We consider $T$-equivariant local coordinates $z_{1},\cdots,z_{N}$ around the image of the $T$-invariant point $p_{\tau}\in U_{\sigma}$ and consider its pullback (as rational functions) to $U_{\tau}\times\mathop{\mathrm{Spec}}\nolimits(R)$ and denote them as $\frac{f_{i}}{g_{i}}$ with $T$-semiinvariant holomorphic (algebraic) $f_{i}$s and $g_{i}$s. Consider the ideal $\mathcal{J}:=\langle\{{\rm lcm}\{g_{i}\},f_{1}\frac{{\rm lcm}\{g_{i}\}}{g_{1}},\cdots,f_{N}\frac{{\rm lcm}\{g_{i}\}}{g_{N}}\}\rangle$ which can be decomposed as
@@
where $\vec{m}$ runs over elements of $\tau^{\vee}\cap M$, $z^{\vec{m}}$ means $\prod_{i}z_{i}^{m_{i}}$, and $I_{\vec{m}}$ are ideals of $R$. With respect to the natural partial order on $\tau^{\vee}\cap M$ as $\vec{m}\leq\vec{m^{\prime}}:\Leftrightarrow\vec{m^{\prime}}-\vec{m}\in\tau^{\vee}\cap M$, we have $I_{\vec{m}}\subset I_{\vec{m}}$ if $\vec{m}\leq\vec{m^{\prime}}$. If we assume $R$ is a discrete valuation ring for simplicity, we take a uniformizer of $R$ as $\pi$ and we can write $I_{\vec{m}}$ as $(\pi^{l_{m}})$, with $l_{m}\in\mathbb{Z}_{\geq 0}$. In this case, we have $l_{\vec{m}}\geq l_{\vec{m^{\prime}}}\geq 0$ if $\vec{m}\leq\vec{m^{\prime}}$.